Gourmet sandwiches. Jenny, Demi, and Drew buy
an 18-in. long gourmet sandwich and take it back to
their apartment. Since they have different appetites,
Jenny cuts the sandwich so that Demi gets half of
what Jenny gets and Drew gets three-fourths of
what Jenny gets. Find the length of each person’s
sandwiches

Answer by CharlesG2(828)

Gourmet sandwiches. Jenny, Demi, and Drew buy
an 18-in. long gourmet sandwich and take it back to
their apartment. Since they have different appetites,
Jenny cuts the sandwich so that Demi gets half of
what Jenny gets and Drew gets three-fourths of
what Jenny gets. Find the length of each person’s
sandwiches
J Jenny
D Demi = 1/2 * J
R Drew = 3/4 * J
J + 1/2 * J + 3/4 * J = 18
4/4 * J + 2/4 * J + 3/4 * J = 18
9/4 * J = 18
9J = 18 * 4 = 72
J = 8 inches
D = 4 inches
R = 6 inches
J + D + R = 8 + 4 + 6 = 12 + 6 = 18 inches